Title

INAF queue observing, 2014 Dec 21-22

Observer: F. Cusano, A. Marchetti
Telescope Operator: G. Bechetti
Support Astronomer: M. Edwards by phone
Instrument: MODS

Summary

Thin cirrus and clouds at sunset. We decide to use MODS with LBC-red, since the sky subtraction with LUCI can be very hard in these conditions.
We stop observations at 2:30UT the sky is completely covered by clouds.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Antoniucci Uz Tau E dual grating   COMPLETED WITH MODS SPEC, imaging missing
Antoniucci V1118 Ori dual grating   COMPLETED WITH MODS SPEC, imaging missing
Antoniucci V1184 Tau dual grating   COMPLETED
00:00 The TO opens the dome

00:36 We take sky flats for LBC-red without dithering
  • filter I-Bessel (good), R-Bessel (not good)

1.32 We run the acquisition for Fusi Pecci U. acqMODS MCEC4_ls.acq (seeing 1.2, thin cirrus)
  • mods1r.20141222.0001-0003

1:48 We run the scientific ob, execMODS MCEC4_ls_10.obs. Thick clouds coming in, guide star almost lost at 2:27UT
  • mods1r.20141222.0004-0005, mods1b.20141222.0001-0002
  • stop integration at 2:30 too many clouds

6:22 The sky looks better, we start again observations

6:33 some clouds at horizon, we try to slew to Saracco_B, acqMODS mods.7.rec.acq
  • mods1r.20141222.0006, seeing 1.5, seeing to bad for this program some clouds in the field (guide star 1.5 mag fainter on sx guideplot)

6:43 slew to Antoniucci Uz Tau E, acqMODS uztaue_ls.acq
  • mods1r.20141222.0006-0009

6:55 execMODS uztaue_ls.obs
  • mods1r.20141222.0010-0013, mods1b.20141222.0003-0006
  • seeing slowing decreasing we try to slew again to Saracco_B

07:14 slewing again to Saracco_B seeing 1.2 on the dimm, acqMODS mods.7.RDCSm1.acq (doing again acquisition after a long slew)
  • mods1r.20141222.0014-0016

07:31 we run the scientific script execMODS mods.7.RDCSm1_4.obs (Saracco_B)
  • mods1r.20141222.0017 seeing jumped to 2 due to atmospheric turbulence. We stop the integration after 1 exposure.

7:52 we slew to Antoniucci v1647 program, acqMODS v1647ori_ls.acq
  • mods1r.20141222.0018-0019 clouds come back all over the sky, the acquisition fails and the TO has to update the pointing, but the object is too faint for the sky conditions

8:00 Marco Stangalini starts performing tests on LBC from Italy

8:06 we slew to Antoniucci v1118ori. acqMODS v1118ori_ls.acq
  • mods1r.20141222.0020-0023

8:19 we run the science execMODS v1118ori_ls.obs
  • mods1r.20141222.0024-0027, mods1r.20141222.0007-0010, seeing reaches 2

8:50 we slew to v1184tau, acqMODS v1184tau_ls.acq
  • mods1r.20141222.0028-0030

9.01 we run the science execMODS v1184tau_ls.obs
  • mods1r.20141222.0031-0034, mods1r.20141222.0011-0014

9:23 execMODS v1184tau.img
  • mods1r.20141222.0035-0037, mods1r.20141222.0015-0017
9:41 weather conditions are getting worse, thick clouds cover 60% of the sky and thin cirrus all around

10:12 we Slew to Saracco_B, acqMODS mods.7.rec.acq, seeing 1.0
  • mods1r.20141222.0038

10:22 execMODS mods.7.RDCSm1_4.obs
  • mods1r.20141222.0039-0042 The guide star gets 1 mag fainter. We stay on target to allow LBC testing

11:38 the TO closes the dome due to humidity

-- %USERSIG{FeliceCusano - 2014-12-19}%

Comments

 
Topic revision: r6 - 23 Jan 2015, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback